# Intelligent Framework for Human-Robot Collaboration: Safety, Dynamic Ergonomics, and Adaptive Decision-Making

### v1

@@ IV-B Experiment: Ergonomic Monitoring, Risk Classification and Robotic Intervention via Behavior Tree
IV-C Experiment: Recognizing Actions with SlowOnly Model
Figure 9: Trend of classification loss during the 200 training epochs. The curve shows stable convergence, suggesting progressive learning of the model.
Figure 10: Increased top-1 accuracy on the validation set during the 200 training epochs. Accuracy shows continuous improvement, highlighting the effectiveness of model optimization..
@@ IV-C Experiment: Recognizing Actions with SlowOnly Model
The framework uses an innovative approach that combines action recognition with ergonomic and volumetric analysis, integrating the results within a Behavior Tree (BT) to determine possible robotic intervention. Video sequences were preprocessed to 8-frame clips, uniformly sampled with a time interval of four frames, and normalized using ImageNet mean values and standard deviations. Each clip was subjected to resizing, center cropping, and augmentation operations with random color changes and horizontal flips to improve the robustness of the model. Training was conducted on an NVIDIA RTX 3090 GPU, using a batch size of 16 videos per GPU.
The optimization process used the AdamW algorithm with an initial learning rate of 0.001, adjusted through a Cosine Annealing strategy with warmup for the first 1,000 steps. Training was performed for 200 epochs, with the cross-entropy loss function. During the validation phase, the model achieved an accuracy of 95,83%, as shown in Fig. 10, demonstrating an excellent ability to generalize over the test data.
The behavior of loss during training shows rapid initial convergence, with stabilization at 0.129 toward the later epochs, as shown in Fig. 10. This result reflects the effectiveness of the model setup and training pipeline in gradually reducing the classification error. The use of short 8-frame clips proved particularly effective in capturing action movements without introducing significant computational complexity, while integration with the BT ensured that the classified data were exploited for real-time robotic decisions.
This model and framework configuration was further validated by the previous evaluation on HRI30, where SlowOnly was shown to outperform alternative architectures such as I3D and TSM in action recognition in industrial settings. The model’s robustness to occlusions and its high accuracy make it an ideal choice for collaborative human-robot scenarios, where accurate and timely action recognition is crucial.
The results of this experiment confirm the effectiveness of the proposed approach, which combines action recognition, ergonomic analysis, and volumetric evaluation, integrating them into a modular and responsive decision-making framework. The framework showed a remarkable ability to adapt to operational variables, providing decisive support for safety and efficiency in collaborative industrial operations.
@@ V Conclusions and Future Work
To our knowledge, the framework represents a novelty in the existing literature, synergistically integrating advanced perception technologies and real-time decision-making. Compared with traditional methods based on static rules or post-hoc ergonomic analysis, our approach introduces greater flexibility and adaptability, improving both the safety and well-being of operators. The noninvasive nature of the system, which eliminates the need for wearable physical sensors, also helps to reduce operational costs and improve the naturalness of human-robot interaction.

### v2

@@ IV-B Experiment: Grasping Action Detection and Volumetric Weight Estimation with Variable Size Parcels
Three segmentation models were analyzed: YOLOv8x-seg, YOLOv9-seg and YOLO11x-seg, trained with the same configuration parameters and evaluated on two main classes, ”parcel” and ”hand”.
For grasping intention recognition, 40 trials were conducted on dynamic sequences, combining the segmentation model with the probabilistic Kalman Unscented (UKF) filter. For volumetric estimation, 4 parcels of varying sizes and shapes were tested, each in 10 trials for a total of 40 measurements. The estimated volumetric weights, obtained from the segmented masks and depth data provided by the RealSense D435 sensor, were compared with the actual weights.
The performance of the segmentation models was evaluated using mAP@50 (mean Average Precision with an IoU threshold of 0.5) and mAP@50:95 (mean Average Precision averaged over multiple IoU thresholds from 0.5 to 0.95 with a step size of 0.05). Precision, defined as the ratio of true positives to the sum of true positives and false positives, and recall, defined as the ratio of true positives to the sum of true positives and false negatives, were calculated for class-specific analysis alongside mAP values.
For volumetric weight estimation, percentage errors were calculated using the formula:
$E_{p}=\frac{|P_{v,\text{st}}-P_{v,\text{re}}|}{P_{v,\text{re}}}\times 100$
where $P_{v,\text{st}}$ e $P_{v,\text{re}}$ represent estimated and actual volumetric weights, respectively.
As shown in Table II, YOLOv8x-seg achieved a mAP@50 of 77.6% and a mAP@50:95 of 71.2%, demonstrating a good balance between precision and recall. YOLOv9-seg slightly improved the mAP@50 (77.7%) and maintained the mAP@50:95 at 72.4%, with more efficient latency management. However, YOLO11x-seg outperformed both, achieving a mAP@50 of 77.8% and a mAP@50:95 of 72.4%, showing greater robustness in complex scenarios.
For grasping intention recognition, the system correctly identified grasping in 37 cases out of 40, with an overall accuracy of 92.5%. The overall mean percent error for volumetric weight estimation over all trials was calculated as 17.58%, with greater variability observed in the extreme size parcels, attributable to the sensor’s limitations in terms of spatial resolution and surface reflectivity.
In conclusion, the model comparison confirmed that YOLO11x-seg represents the best performing solution for the operational context considered. Its combination of robustness, accuracy, and ability to handle difficult conditions, such as detection of moving or partially occluded hands, makes it ideal for grasping and volumetric estimation applications. The experimental outcomes reveal the advantages of integrating a deep learning model with probabilistic tracking algorithms, demonstrating superior robustness compared to the isolated use of neural networks. However, volumetric weight estimation showed sensitivity to the quality of depth data provided by the sensor, with a reduction in accuracy due to limitations in spatial resolution or surface reflectivity. This suggests that further improvements could be achieved through refinement of sensor calibration or integration of data from multiple sensors to compensate for current limitations.
@@
IV-C Experiment: Ergonomic Monitoring, Risk Classification and Robotic Intervention via Behaviour Tree
The objective of this experiment is to evaluate the framework’s ability to monitor operator postures in real time, classify ergonomic risk, and manage task transfer to the robot in the presence of critical conditions. During testing, the operator performed movements typical of an industrial environment, such as bending and lifting, while the system monitored posture using OpenPose and classified ergonomic risk using the OWAS method.
When medium-high-risk (OWAS class 3) or high-risk (OWAS class 4) postures were detected, or volumetric weight in excess of allowable limits, the system generated a biomechanical fatigue message. This message was processed by a BT, which combined ergonomic data, action recognition results, and volumetric weight values to determine whether robotic intervention should be activated. This approach ensured modular and responsive decision making, allowing a smooth transition between the operator and the robot.
The performance of this experiment was evaluated through several metrics: the distribution of postures across OWAS risk classes (1-4), system update rate (Hz) and latency (seconds) for pose monitoring, response time (seconds) to risky situations, and time trend analysis of posture risk levels.
The data collected during the experiment showed that 12,8% of the postures belonged to OWAS class 1 (no risk), 84,6% belonged to OWAS class 2 (moderate risk), no postures were classified in OWAS class 3 (medium-high risk), and 2,6% belonged to OWAS class 4 (high risk), as shown in Fig. 6. These results confirm the system’s ability to distinguish between ergonomically safe and hazardous postures, demonstrating the effectiveness of the framework in monitoring and classifying risk. Compared to the approach based on OpenPose by Lin et al. [45], our system demonstrates significantly higher precision in ergonomic evaluation. The classification of high-risk postures (OWAS class 4) stands at 2.6%, drastically reducing false positives compared to the 10.4% reported in the comparative work. This enhanced diagnostic precision minimizes unnecessary operational interruptions, a crucial aspect in dynamic industrial environments.
Temporally, the system maintained an average update rate of 14.99 Hz, with an average latency of 0.57 seconds for pose monitoring, as shown in Fig. 6. These values ensured that any postural changes were detected and analyzed in a timely manner. In addition, the average response time to risky situations was 0.07 seconds, as shown in Fig. 8, demonstrating the efficiency of the framework in identifying critical conditions and responding quickly. The system’s reactivity, measured through the average response time to ergonomic risk situations, reaches 0.07 seconds, exceeding by more than 56% the benchmark of 0.16 seconds reported by Tortora et al. [46]. This substantial improvement in latency was achieved while maintaining a recognition accuracy of 92.5%, comparable to the 94.3% of the reference system. The optimal balance between speed and precision enables our framework to implement real-time intervention capabilities previously unattainable in dynamic industrial environments, where even a few milliseconds can determine the effectiveness of safety measures.
These measurements demonstrate not only the framework’s efficiency in detecting and analyzing critical conditions, but also its role in supporting a seamless operational transition. The reduced overall latency, as calculated through pose monitoring and response time to risky situations, ensured that the transfer of activities to the robot was smooth and responsive. This level of operational readiness, combined with the system’s ability to integrate ergonomic and actionable data, represents a significant step toward optimizing human-robot collaboration.
IV-D Experiment: Recognizing Actions with SlowOnly Model
@@
Video sequences were preprocessed to 8-frame clips, uniformly sampled with a time interval of four frames, and normalized using ImageNet mean values and standard deviations. Each clip was subjected to resizing, center cropping, and augmentation operations with random color changes and horizontal flips to improve the robustness of the model. The optimization process used the AdamW algorithm with an initial learning rate of 0.001, adjusted through a Cosine Annealing strategy with warmup for the first 1,000 steps. Training was conducted on an NVIDIA RTX 3090 GPU, using a batch size of 16 videos per GPU for 200 epochs, with the cross-entropy loss function.
The experiment’s performance was assessed through classification loss during training, top-1 accuracy on the validation set, comparison with alternative architectures (I3D and TSM), and model robustness to occlusions and variability.
During the validation phase, the model achieved an accuracy of 95,83% demonstrating an excellent ability to generalize over the test data. The behaviour of loss during training shows rapid initial convergence, with stabilization at 0.129 toward the later epochs. This result reflects the effectiveness of the model setup and training pipeline in gradually reducing the classification error.
The use of short 8-frame clips proved particularly effective in capturing action movements without introducing significant computational complexity, while integration with the BT ensured that the classified data were exploited for real-time robotic decisions.
This model and framework configuration was further validated by the previous evaluation on HRI30, where SlowOnly was shown to outperform alternative architectures such as I3D and TSM in action recognition in industrial settings. Comparative analysis [22] revealed that SlowOnly achieved a top-1 accuracy of 86.55% compared to 78.43% for I3D and 73.91% for TSM when evaluated on the HRI30 dataset. The model’s robustness to occlusions and its high accuracy make it an ideal choice for collaborative human-robot scenarios, where accurate and timely action recognition is crucial.
Experimental outcomes establish the effectiveness of the proposed approach, which combines action recognition, ergonomic analysis, and volumetric evaluation, integrating them into a modular and responsive decision-making framework. The framework showed a remarkable ability to adapt to operational variables, providing decisive support for safety and efficiency in collaborative industrial operations.
@@ V Conclusions and Future Work
To the best of our knowledge, this is the first framework to synergistically integrate advanced visual perception, real-time ergonomic assessment, and adaptive decision-making for industrial HRC. Compared with traditional methods based on static rules or post-hoc ergonomic analysis, our approach introduces greater flexibility and adaptability, improving both the safety and well-being of operators. The noninvasive nature of the system, which eliminates the need for wearable physical sensors, also helps to reduce operational costs and improve the naturalness of human-robot interaction.
@@
  • Cross-domain Applications: Explore the framework’s adaptability in non-industrial contexts such as healthcare (surgical assistance and patient mobility support), smart homes (assistive living for elderly individuals), rehabilitation centers (personalized therapy tracking), and logistics (warehouse automation and order fulfillment).